# Bourbaki degree of pairs of projective surfaces

### v2

@@ 5 $e\leq 1$
Here the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the codimension two part of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of a line $V(x_{1},x_{2})$ with multiplicity two. The saturation of the first Fitting ideal is $(0)$ in this case. The sheaf ${\mathcal{T}}_{\sigma}$ admits the following free resolution:
@@
so that $e=2$, $\operatorname{Bour}(\sigma)=3$ and $m(\sigma)=2$, with $c_{3}({\mathcal{T}}_{\sigma})=7$.
@@
Here, the saturation of the 0-th Fitting ideal coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $(1)$. Moreover, $m(\sigma)=2$ with $\operatorname{supp}(\mathcal{Q}_{\sigma})$ being a double line structure at $V(x_{0},x_{3})$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
@@
Item $(a)$ is 46 and item $(b)$ follows with 51. Item $(c)$ is shown in 53, and the stability result in $(d)$ is in 58. ∎

### v3

@@ 5 $e\leq 1$
The irreducible components of $(\Xi_{\sigma})_{1}$ consist only of a line $V(x_{1},x_{2})$. Consider $\mathfrak{p}=(x_{1},x_{2})$ the corresponding prime ideal in $R$, and let $u=(x_{2}-x_{0})^{-1}\in R_{\mathfrak{p}}$. The elementary operations $R_{2}^{\prime}=R_{2}-2ux_{0}x_{1}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ turn the matrix into
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&1&u(x_{1}-x_{3})&-ux_{2}\\ x_{1}^{2}+2ux_{0}x_{1}^{2}&0&x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})&x_{2}^{2}+2ux_{0}x_{1}x_{2}\end{pmatrix}$,
which sends $e_{2}\in R_{\mathfrak{p}}^{4}$ to $(1,0)\in R_{\mathfrak{p}}^{2}$, so we may compute the cokernel $(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ by considering the second row and the remaining entries:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I},$
where
$f_{1}$ $=x_{1}^{2}(1+2ux_{0})$
$f_{2}$ $=x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})$
$f_{3}$ $=x_{2}(x_{2}+2ux_{0}x_{1}).$
Since $1+2ux_{0}\in R_{\mathfrak{p}}^{\times}$ and $f_{1}\in I$, we conclude $x_{1}^{2}\in I$. This simplifies $f_{2}\in I$ to:
$x_{2}(3x_{2}+2x_{3})\equiv-(2ux_{0}x_{3})x_{1}\mod I,$
so there is a unit $v\in R_{\mathfrak{p}}^{\times}$ such that $x_{1}\equiv vx_{2}\mod I$, which together with $x_{1}^{2}\in I$ implies $x_{2}^{2}\in I$. From this, $f_{3}\in I$ simplifies to
$f\equiv 2ux_{0}x_{1}x_{2}\in I\Rightarrow x_{1}x_{2}\in I,$
and therefore we have
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{1}\rangle,$
with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=2$. Thus, $m(\sigma)=2$ (from the associativity formula), and $e=2$, yielding $\operatorname{Bour}(\sigma)=3$. The following resolution may also be obtained computationally:
@@
describing the Bourbaki scheme as a twisted cubic and giving $c_{3}({\mathcal{T}}_{\sigma})=7$.
@@
Here, there is a unique irreducible component of $(\Xi_{\sigma})_{1}$, the line $V(x_{0},x_{3})$. Considering $\mathfrak{p}=(x_{0},x_{3})$ the associated prime ideal, we compute $m(\sigma)=2$ using the associativity formula.
Take $u=(x_{2})^{-1}\in R_{\mathfrak{p}}$ and consider the elementary operations $R_{2}^{\prime}=R_{2}-u(x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2})R_{1}\doteq R_{2}-uQR_{1}$, $R_{1}^{\prime}=uR_{1}$ so that
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&-ux_{0}&ux_{3}&1\\ 2x_{0}x_{2}+x_{1}x_{3}+uQx_{1}&x_{0}x_{3}+uQx_{0}&x_{0}^{2}+x_{3}^{2}-uQx_{3}&0\end{pmatrix}$.
Arguing as usual, we may compute the cokernel by
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I}$
with
$f_{1}$ $=x_{0}(2x_{2}+ux_{1}^{2})+x_{1}x_{3}(2ux_{2}+3ux_{3}+1)$
$f_{2}$ $=x_{0}(x_{3}(2ux_{2}+3ux_{3}+1)+ux_{0}x_{1})$
$f_{3}$ $=x_{0}^{2}+x_{3}(x_{3}-ux_{0}x_{1}-2ux_{2}x_{3}-3ux_{3}^{2})$
Writing
$f_{2}-\frac{x_{0}}{x_{1}}f_{1}$ $=x_{0}\left(-\frac{x_{0}}{x_{1}}(2x_{2}+ux_{1}^{2})+ux_{0}x_{1}\right)$
$=x_{0}^{2}\left(-\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)\in I,$
and since the term inside parenthesis is invertible in $R_{\mathfrak{p}}$, we conclude $x_{0}^{2}\in I$.
On the other hand, we consider
$B\doteq f_{3}-x_{0}^{2}+\frac{x_{3}}{x_{1}}f_{1}$ $=-x_{3}\left(x_{0}\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{0}x_{1}-2x_{3}\right)$
$=-x_{3}x_{0}\left(\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)+2x_{3}^{2}\in I.$
and moreover, let
$A\doteq x_{0}x_{3}(2ux_{2}+3ux_{3}+1)=f_{2}-ux_{0}^{2}x_{1}x_{3}\in I.$
Then
$A-\frac{3}{2}ux_{0}B$ $=x_{0}x_{3}(2ux_{2}+1)+3x_{3}^{2}-3x_{3}^{2}+\frac{3}{2}ux_{0}^{2}x_{3}\left(\frac{-(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right),$
and after removing the term with $x_{0}^{2}$ factor, we obtain
$x_{0}x_{3}(2ux_{2}+1)\in I\Rightarrow x_{0}x_{3}\in I,$
since $2ux_{2}+1$ is invertible in $R_{\mathfrak{p}}$. Now, removing the $x_{0}x_{3}$ term in $B$ we conclude $x_{3}^{2}\in I$. So, up until now, we have $x_{0}^{2},x_{0}x_{3}$ and $x_{3}^{2}\in I$. Finally, considering
$f_{1}-3ux_{1}x_{3}^{2}=x_{0}(2x_{2}+ux_{1}^{2})+x_{3}x_{1}(2ux_{2}+1)\in I,$
this means there are $a,b\in R_{\mathfrak{p}}^{\times}$ such that $ax_{0}+bx_{3}\in I$. Therefore, we conclude that
$I=(x_{0}^{2},x_{3}^{2},x_{0}x_{3},ax_{0}+bx_{3}),$
and these mean that we can write
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{0}\rangle\simeq\langle 1,x_{3}\rangle$
and obtain that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=1$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
Theorem D.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$. Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
  • (e)
    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
@@
Item $(a)$ is 48 and item $(b)$ follows with 53. Item $(c)$ is shown in $5$, and item $(d)$ is described in 57. The stability result in $(e)$ is in 60. ∎